### v1

# Meta-Reasoner: Dynamic Guidance for Optimized Inference-time Reasoning in Large Language Models

Abstract
Large Language Models (LLMs) increasingly rely on prolonged reasoning chains to solve complex tasks. However, this trial-and-error approach often leads to high computational overhead and error propagation, where early mistakes can derail subsequent steps. To address these issues, we introduce Meta-Reasoner, a framework that dynamically optimizes inference-time reasoning by enabling LLMs to “think about how to think.” Drawing inspiration from human meta-cognition and dual-process theory, Meta-Reasoner operates as a strategic advisor, decoupling high-level guidance from step-by-step generation. It employs contextual multi-armed bandits to iteratively evaluate reasoning progress, and select optimal strategies (e.g., backtrack, clarify ambiguity, restart from scratch, or propose alternative approaches), and reallocates computational resources toward the most promising paths. Our evaluations on mathematical reasoning and puzzles highlight the potential of dynamic reasoning chains to overcome inherent challenges in the LLM reasoning process and also show promise in broader applications, offering a scalable and adaptable solution for reasoning-intensive tasks.
Meta-Reasoner: Dynamic Guidance for Optimized Inference-time
Reasoning in Large Language Models
Yuan Sui1, Yufei He1, Tri Cao1, Simeng Han2, Bryan Hooi1 1 National University of Singapore 2 Yale University {yuansui, bhooi}@comp.nus.edu.sg
1 Introduction
o1-like reasoning chains allow Large Language Models (LLMs) to “think for an extended period” before actually solving a problem. This shows impressive performance on challenging tasks, such as logical problems puzzles Lei et al. (2024); Yao et al. (2023), math questions Patel et al. (2024); Lightman et al. (2023), logical reasoning Han et al. (2024), and science questions Rein et al. (2023), which often pose difficulties for even the most advanced models Gandhi et al. (2024); Sui et al. (2024c); He et al. (2024).
However, the trial-and-error nature of o1-like reasoning often incurs substantial computational overhead Snell et al. (2024); Manvi et al. (2024) and is prone to error propagation where early flaws in a reasoning chain can compound and derail subsequent steps Lei et al. (2024); Yao et al. (2023); Gandhi et al. (2024). While related iterative approaches Gandhi et al. (2024); Li et al. (2025) have explored techniques like partial revision or backtracking, they typically address errors in an ad-hoc manner for a narrow span of reasoning steps and lack a systematic way to assess whether an entire line of reasoning remains viable. Thus, models remain vulnerable to getting “stuck” in less promising reasoning trajectories, continuously expending computational resources on unpromising paths rather than recognizing when a major strategic shift is needed. A critical challenge, therefore, is to enable LLMs to allocate their reasoning budget more effectively, prioritizing promising avenues while adapting—or discarding—ineffective strategies.
Figure 1: High-level comparison of LRM with o1-like chains and meta-reasoner.
To overcome these challenges, we propose Meta-Reasoner, a specialized module that operates alongside the LLM to enhance its reasoning capabilities. The meta-reasoner serves as an “advisor”, dynamically evaluates the reasoning process, offering high-level guidance and strategic redirection when progress stalls. Unlike the LLM, which focuses on more specific stepwise generation, the meta-reasoner focuses on a broader perspective and evaluates the overall progress and strategy of the reasoning process. Meta-Reasoner operates in iterative rounds: where the LLM first generates partial o1-like reasoning chains and provides a “progress report” summarizing its reasoning progress so far. Based on these reports, the meta-reasoner then provides the strategic guidance, such as restarting with a different approach, refining existing ideas, or targeting specific sub-problems. Crucially, the meta-reasoner is designed to not interact with the granular details of the CoT; instead, it focus on the global oversight of the reasoning progress and provides dynamic high-level strategies, preventing the LLM from getting stuck or wasting resources on unproductive lines of inquiry.
Overall, the main contributions of this paper are summarized as follows:
  • We propose a novel meta-reasoning framework that operates as a high-level advisor for LLMs, enabling them to “think about how to think” by dynamically optimizing inference-time reasoning strategies.
  • By decoupling global strategy decisions from low-level chain-of-thought generation, Meta-Reasoner oversees progress through concise “progress reports” rather than micromanaging each reasoning step. This design mitigates error propagation and reduces wasted computation on unproductive paths.
  • We evaluate Meta-Reasoner on challenging mathematical and scientific reasoning benchmarks (e.g., Game of 24, TheoremQA, and SciBench), demonstrating significant improvements in both accuracy and efficiency compared to baselines. Our results show that the proposed framework offers a scalable solution to inference-time reasoning bottlenecks.
2 Related Works
Complex Reasoning in LLMs
The introduction of CoT reasoning has revolutionized how LLMs approach problem-solving, allowing them to break tasks into intermediate steps (Lee et al., 2025). The recent LLMs like o1, o3 from OpenAI and Deepseek-v3 from Deepseek have achieved state-of-the-art results in diverse domains using CoT-like reasoning (Manvi et al., 2024; Li et al., 2025; Kudo et al., 2024; Sui et al., 2024b). However, CoT’s sequential dependency limits its robustness, as errors in earlier steps can cascade through the process (Snell et al., 2024) and also when facing complex reasoning tasks (Sui et al., 2024a, c), CoT-like reasoning may stuck in the infinite loop of reasoning (Lee et al., 2025). These issues motivate us to propose Meta-Reasoner to assess and adapt the overall reasoning strategy based on the progress of CoT reasoning. Unlike the LLMs, which focuses on more specific each step generation, the meta-reasoner focus on the border perspective and evaluates the overall progress and strategy of the reasoning process. It can provide a global oversight to avoid LLMs getting stuck or wasting resources on unproductive lines of thoughts.
Backtracking and Error Correction
Addressing cascading errors in multi-step reasoning remains a central challenge. Recent approaches have focused on backtracking and self-verification Yao et al. (2023); Besta et al. (2023); Gandhi et al. (2024). For instance, Weng et al. (2023) have shown that incorporating a self-verification stage—where the model re-checks its conclusions using the very chain of thought it generated—can dramatically boost performance by catching missteps early. Similarly, Ling et al. (2023) propose not only generate multiple candidate reasoning chains but also employs a verifier mechanism to identify and backtrack on erroneous steps. These techniques go beyond post-hoc validation by introducing dynamic strategy adjustments during inference Lightman et al. (2023), thereby reducing the impact of errors propagating through long reasoning chains. Following these useful efforts, we initiate our Meta-Reasoner with the instructions like (1) restart from scratch and propose alternative strategies; (2) backtracking to the point where the error occurred; and (3) continue and provide specific suggestions. The detailed strategy can be found in §4.3.
Meta-Cognition & Dual-Process Systems
Meta-cognition in human reasoning involves higher-order processes that allow individuals to monitor, evaluate, and adjust their cognitive strategies (Gao et al., 2024; Yoran et al., 2024). This reflective thinking—often seen as System 2 processes in dual-process theories Havrilla et al. (2024)—is vital for tasks that require careful deliberation and error correction Didolkar et al. (2024). Drawing on these insights, our Meta-Reasoner can be considered analogous to dual-process systems, where LRM for generating CoT steps parallels System 1 and Meta-Reasoner for providing high-level strategic oversight to guide or redirect reasoning when needed serves as System 2. This separation of responsibilities enables the framework to balance efficiency with robust problem-solving, where the LRM handles routine inferences and the meta-reasoner intervenes when high-level strategic adjustments.
3 Preliminary
A central challenge in complex reasoning tasks is choosing the most effective strategy among multiple valid options. This challenge naturally aligns with the contextual multi-armed bandit (MAB) framework, which is designed to balance the exploration of new strategies with the exploitation of strategies known to perform well.
In this framework, an agent observes a context $x_{t}$ that characterizes the current state of the environment at every time step $t$ and selects an arm $s_{t}$ from a finite set $\mathcal{S}$. Upon selecting arm $s_{t}$, the agent receives a reward $r(s_{t},x_{t})$ that reflects both the chosen arm and the context. The primary objective of MAB is to maximize the cumulative reward over time: $R(T)=\sum_{t=1}^{T}r(s_{t},x_{t})$. A central challenge in the contextual MAB problem is balancing exploration (trying different arms to gather information about their rewards) with exploitation (selecting the arm that has yielded high rewards in similar contexts in the past). This balance ensures that while the agent leverages known profitable actions, it also continues to search for potentially better options. This principle is central to our motivation behind Meta-Reasoner, which aims to automatically select the most efficient strategy to guide the reasoning process during inference time.
A widely used algorithm in the contextual setting is LinUCB Li et al. (2012), which models the expected reward as a linear function of the context. Specifically, for an arm $s$ given context $x_{t}$, the expected reward is modeled as $\mathbb{E}\left[r\left(s,x_{t}\right)\right]\approx x_{t}^{\top}\theta_{s}$, where $\theta_{s}$ is an unknown parameter vector associated with arm $s$. To manage uncertainty in its estimates, LinUCB maintains for each arm an estimate $\hat{\theta}_{s}$ and an associated covariance matrix $A_{s}$. The algorithm then selects the arm according to:
$s_{t}=\arg\max_{s\in\mathcal{S}}\left[x_{t}^{\top}\hat{\theta}_{s}+c\sqrt{x_{t}^{\top}A_{s}^{-1}x_{t}}\right],$ (1)
where $c$ is a constant that controls the exploration level. Here, the term $c\sqrt{x_{t}^{\top}A_{s}^{-1}x_{t}}$ serves as a confidence bound on the reward estimate, encouraging the selection of arms with higher uncertainty (i.e., those with less historical data) and thereby facilitating exploration. By incorporating context into the decision-making process, LinUCB allows the agent to adapt its strategy based on the current state, aligning well with our goal of selecting the most efficient reasoning strategy under varying conditions.
4 Methods
Based on the intuition to allow the LLMs to focus on their computation on more promising lines, we are motivated by two research questions: (1) How can we enable language models to dynamically allocate resources during inference to optimize for reasoning and planning? (2) What architecture allows for effective separation between the reasoning process in LRM and the meta-level guidance of that process in Meta-reasoner? To address them, we propose a new framework, Meta-Reasoner, to equip the LLMs with the capabilities to “thinking about how to think”. It supervises the reasoning process of the LLMs and provides dynamic strategies to enable the LLMs to focus on more promising reasoning paths instead of iterative “trial-and-error”. This framework also addresses limitations of the current sequential generation of the reasoning paths which may get stuck in suboptimal trajectories by balancing “higher-order” thinking.
Figure 2: An illustration of the Meta-Reasoner workflow. In each round, the LLM produces a new reasoning step to extend its CoT reasoning. The CoT is then summarized into a progress report, which provides context for the meta-reasoner. Then meta-reasoner uses a contextual multi-armed bandit (either using a fixed contextual bandit or dynamic contextual bandit) to choose a guidance strategy. The selected strategy then guides the next reasoning step generation, to enable strategic redirection, error correction, and resource optimization. A reward is then computed from the progress report and used to update the bandit algorithm. The process repeats until the task is complete or the maximum number of rounds is reached.
The meta-reasoning framework operates iteratively as shown in Figure 2. At each round $t$, the reasoning process comprises three steps: (1) CoT generation by the LLM, (2) Progress Reporting to summarize the reasoning progress so far (i.e., this is partly for efficiency, and partly to help the meta-reasoner focus on its main goal of “advising” rather than being distracted by the details in the CoT), and (3) Strategy Generation by the meta-reasoner to optimize subsequent steps. The selection of the strategy is almost exactly corresponds to the well-studied problem of contextual multi-armed bandits learning. Each strategy can be seen as an arm for the bandit, and the reward of each strategy can be evaluated by the progress of LLM reasoning after applying the strategy. We analogy the process of executing and evaluating each strategy as the act of “pulling” each arm. The overall goal of our meta-reasoner is to find the best arm (i.e., strategy with highest cumulative rewards) with as few pulls as possible. The complete algorithm of Meta-Reasoner is appended in Algorithm 1.
4.1 Chain-of-Thought (CoT) Generation
In the first step, the LLM generates a reasoning step to extend its CoT reasoning based on the user query. Starting with its reasoning history $C_{t-1}$ and the guidance $G_{t-1}$ provided by the meta-reasoner in the previous round, the LRM $M$ produces a new reasoning step $s_{t}$. This step is then appended to the current CoT, forming $C_{t}=C_{t-1}\cup\{s_{t}\}$. This incremental process allows the LRM to iteratively build a structured reasoning path. By keeping track of the full reasoning trajectory at each round, the model creates a coherent foundation for evaluation and further refinement. This process is alike the demonstration in o1-like models, which generate a long-term thinking process. However, the issue of this reasoning is that its more like a process of “trial-and-error”, which may waste some of the inference costs on unnecessary/useless paths. In addition, due to the sequential generation process, it may easily get stuck in suboptimal solutions.
4.2 Progress Reporting
Once the LRM has updated its CoT, we summarize the reasoning history $C_{t}$ into a concise progress report $P_{t}$. This summary captures the key aspects of the reasoning trajectory, such as how much progress has been made toward the task goal, the consistency of the reasoning, and any significant updates so far. The summarization function $f$ abstracts the detailed CoT into a simpler, more focused representation. This step is designed to be both computationally efficient and informative, ensuring that the meta-reasoner can focus on evaluating high-level progress without being overwhelmed by the granular details of every reasoning step. Even this step is more like an engineering trick, but we find that it may unlock some of the capabilities of LRM to do “higher-order” thinking, we find that with more essential information included in the prompt, the LRM will generally produce more insightful, critical strategies which could be more useful for complex reasoning.
Diagnosis Strategy
Progress is insufficient or the current strategy seems ineffective. Restart from scratch and propose alternative strategies.
[…]
There are mistakes in intermediate steps. Backtrack to the point where the error occurred.
[…]
The current approach is working well. Continue and provide specific suggestions for the next steps.
Ambiguous or conflicting intermediate results are observed. Pause to clarify and disambiguate the current reasoning, then reconcile the discrepancies.
[…]
The reasoning process appears overly complex or convoluted. Simplify by decomposing the task into smaller, manageable sub-tasks.
[…]
Evidence of error propagation or low confidence in certain subcomponents. Perform targeted verification on critical steps and focus on areas with low confidence.
[…]
Repetitive or circular reasoning patterns are detected. Reset to a previously successful checkpoint and explore alternative solution paths.
Table 1: Demonstration: Contextual bandit pair (i.e., diagnosis of the current state and corresponding strategy) for guiding the LLM’s reasoning process. Marked rows are some of the unique strategies generated by Dynamic Contextual Bandits.
4.3 Meta-reasoner Strategy Generation
In the next step, the meta-reasoner evaluates the progress report $P_{t}$ and selects proper strategy $G_{t}$ for LLM reasoning (the complete process can be found in Algorithm 1). We formulate the generation of strategy as a multi-armed bandits problem and consider two settings below: (1) our approach begins with a fixed-strategy formulation, where the meta-reasoner selects from a predefined set of strategies using a contextual bandit algorithm. We then extend this architecture to (2) an advanced setting in which the meta-reasoner is itself an LLM-based agent and can introduce or refine new strategies on the fly. In both cases, the meta-reasoner uses the same partial-feedback principle of multi-armed bandits to adaptively choose which strategy to deploy based on a reward function. The reward function evaluates the quality of the given reasoning progress after applying the meta-reasoner strategy. We demonstrate the contextual bandit pair (i.e., diagnosis of the current state from the progress report and the corresponding strategy) in Table 1.
Progress Evaluation.
A central goal of our evaluation mechanism is to measure how effectively the model’s current reasoning is advancing toward the task objective (e.g., solving a complex problem) while also monitoring computational expenditure to encourage efficiency. Concretely, we implement a reward function that tracks both solution progress (e.g., partial correctness, compliance with constraints) and resource usage (e.g., the number of reasoning steps). In principle, any suitable evaluator can be employed, including LLM-based verification or external scoring scripts. In our setup, we leverage an LLM as evaluator (with prompt referred at Figure 58) to verify the reasoning progress. We adjust the implementation to output a cumulative score which will be further leveraged to update the MAB algorithms.
Fixed Contextual Bandit.
In the basic version of our framework, the meta-reasoner is modeled as a single contextual bandit that selects from a fixed, finite set of $K$ strategies. These strategies may include instructions such as “continue and provide specific suggestions”, “restart from scratch”, “backtrack to the point where the error occurred”, or “propose alternative methods or perspectives to consider”. At each round, the LRM produces a progress report summarizing its partial reasoning, the meta-reasoner transforms this progress report into a feature vector $x_{t}$ using a language model and applies a contextual bandit algorithm (e.g., LinUCB Li et al. (2012)) to select the best next strategy $a_{t}$. The LLM then executes that strategy and we collect the reward $r_{t}$ for $a_{t}$ based on the reward function. Through iterative MAB algorithm updating, the model learns to select a proper strategy based on the context from recent progress report.
Dynamic Contextual Bandit.
The basic framework assumes a static set of arms (strategies). In practice, the meta-reasoner may also be an LLM, capable of inventing new approaches over time. To accommodate dynamic strategies, we allow the meta-reasoner to propose or refine new strategies at round $t$, which generates an expanding collection of actions, $\mathcal{A}_{1}\subseteq\cdots\subseteq\mathcal{A}_{t}$. Each newly proposed strategy becomes an arm in the contextual bandit. To encourage at least some exploration on this new arm, we initialize each arm with a blank or weak prior in bandit’s parameters.
By explicitly separating low-level content generation (handled by the LLM) from high-level strategy decisions (governed by the meta-reasoner’s bandit), the system can effectively avoid getting stuck or wasting excessive resources on poor solution paths. In domains where a predefined set of strategies is sufficient, the fixed-arm formulation can simplify the method deployment. While in more open-ended domains where novel tactics may emerge, dynamic-arm extensions give meta-reasoner more freedom to evolve.
5 Experiments
In this section, we first introduce the experiment settings including the dataset, baselines, and the backbone models. We then present the main results of Meta-Reasoner with some other analysis from perspectives like efficiency, rewards accumulation, and qualitatively assess meta-reasoner output.
5.1 Experiments Setup
Datasets.
We consider the tasks requiring complex reasoning and the proper solutions naturally composed of long reasoning steps. We evaluate Meta-Reasoner on several challenging datasets: the 24-point game proposed by Yao et al. (2023), college-level scientific problem from SciBench Wang et al. (2024) and theorem-driven math question in TheoremQA Chen et al. (2023). For the SciBench, we only consider the math-related subset for testing which covers the diff, stat, and calc (the detailed clarification of each subset collection can be found in Wang et al. (2024) and we provide the demonstration for each subset in Figure 9). For the TheormQA, we consider the mathematics subset that involves logical reasoning for our testing. We follow the experimental setup in MACM Lei et al. (2024) to conduct the corresponding analysis on these datasets.
Baselines.
We consider several established prompting methods as baselines as follows:
  • Chain-of-thought (CoT) Wei et al. (2022): A prompting technique that encourages models to generate intermediate reasoning steps to enhance problem-solving capabilities.
  • Self-Consistent Chain of Thought (SC-CoT) Wang et al. (2022): An extension of CoT that improves reasoning consistency by generating multiple reasoning chains and selecting the most consistent answer.
  • Multi-Chain Reasoning (MCR) Yoran et al. (2024): enhances SC-CoT by having another LLM to assess and integrate content among the sampled reasoning chains to generate the final consistent answer.
  • Tree of Thoughts (ToT) Yao et al. (2023): A method that explores multiple reasoning paths in a tree structure, allowing the model to consider various possibilities before arriving at a conclusion by tree search algorithms.
  • Reflexion Shinn et al. (2024): A framework that enables models to reflect on their reasoning process, iteratively refining their answers based on feedback.
  • MACM Lei et al. (2024): A multi-agent system to refine the reasoning based on iterative condition mining.
Method Diff(%) Stat(%) Calc(%)
Phi-4 + CoT 17.42 28.42 32.93
Llama-3.1-instruct + CoT 33.14 49.72 54.18
Gemini-Exp-1206 + CoT 36.32 56.73 59.24
Gemini-Exp-1206 + SC-CoT 38.73 59.12 64.11
GPT-4o-mini + CoT 33.12 55.71 58.10
GPT-4o-mini + SC-CoT 37.33 56.67 63.81
GPT-4o-mini + MCR 40.12 58.21 67.42
[…]
GPT-4o-mini + MACM Lei et al. (2024) 54.78 67.13 65.77
GPT-4o + MACM Lei et al. (2024) 61.42 78.32 76.72
GPT-4o-mini + Meta-Reasoner (our work) 60.32 73.64 80.23
GPT-4o + Meta-Reasoner (our work) 67.14 83.29 84.17
Table 2: Accuracy (%) comparison of different methods on the math-related subset of the SciBench dataset. Each column refers to the problem subset defined in Wang et al. (2024).
Method Accuracy (%)
GPT-4o-mini + CoT Yao et al. (2023) 4
GPT-4o-mini + SC-CoT Yao et al. (2023) 9
GPT-4o-mini + IO (best of 100) Yao et al. (2023) 33
GPT-4o-mini + CoT (best of 100) Yao et al. (2023) 49
Gemini-Exp-1206 + IO (best of 100) Yao et al. (2023) 38
Gemini-Exp-1206 + CoT (best of 100) Yao et al. (2023) 60
[…]
GPT-4o-mini + ToT ($b=1$Yao et al. (2023) 45
GPT-4o-mini+ ToT ($b=5$Yao et al. (2023) 74
GPT-4o-mini + Reflexion Shinn et al. (2024) 53
GPT-4o-mini + MACM Lei et al. (2024) 80
GPT-4o-mini + Meta-Reasoner (our work) 89
GPT-4o + Meta-Reasoner (our work) 92
Gemini-Exp-1206 + Meta-Reasoner (our work) 94
[…]
o1-mini + IO 89
o1-preview + IO 93
Table 3: Accuracy(%) comparison of different prompting methods on 24-points game. $b$: Search breadth.
Method Accuracy (%)
GPT-4o-mini + CoT 39.46
Gemini-Exp-1206 + CoT 43.12
[…]
GPT-4o-mini + Reflexion Shinn et al. (2024) 74.32
GPT-4 Turbo + MACM Lei et al. (2024) 79.41
GPT-4o-mini + Meta-Reasoner (our work) 84.13
Gemini-Exp-1206 + Meta-Reasoner (our work) 86.32
Table 4: Accuracy(%) comparison of different prompting methods on TheoremQA Chen et al. (2023).
Backbone Models.
We consider both LLMs and the recent LRMs for our experiments. For the LLMs, we consider the closed-source models like GPT-4o, GPT-4o-mini (between Nov 2025 to Jan 2025) from OpenAI, and open-sourced models like meta-llama-3.1-8B-instruct from Meta, phi-4 from Microsoft and gemini-experimental-1206 from Google. For the LRMs, we consider the closed-source models like o1, o1-mini (In case we cannot break down the generation of o1 models through APIs, we cannot properly inject our meta-reasoner with o1-series models; we only provide the IO results for references), and open-sourced models like QwQ-32B-preview from QWen and Deepseek-v3 from Deepseek-AI. For the feature extraction mentioned in §4.3, we use text-embedding-3-small from OpenAI as the embedding model.
To ensure the reproducibility of the experiments, we set $\mathrm{temperature}=0.7$ and $\mathrm{top\_p}=1.0$ for all models. We use the API service from OpenAI [^1] https://openai.com/ and OpenRouter [^2] https://openrouter.ai/ for our experiments which host detailed snapshots of the utilized model versions.
5.2 Main Results
We compare the accuracy of different prompting methods across different backbone models on SciBench (as shown in Table 2), 24-points game (as shown in Table 3) and TheoremQA (as shown in Table 4). We found that basic prompting strategies, such as CoT and SC-CoT, show limited effectiveness, achieving only 4% and 9% accuracy on 24-point games, respectively. Incorporating IO strategy with “Best of 100” samples improves accuracy to 33%, but it remains far behind advanced methods. Strategies like ToT illustrate the importance of exploring broader reasoning paths, with accuracy increasing from 45% to 74% as the search breadth expands from 1 to 5. Advanced iterative methods, such as Reflexion (53%) and MACM (80%), further demonstrate the value of refined reasoning frameworks. Our proposed Meta-Reasoner outperforms these approaches, achieving 89% accuracy with GPT-4o-mini and 92% with GPT-4o, showcasing its ability to dynamically guide reasoning, correct errors, and focus resources effectively. Compared to specialized models like o1-mini, our method equipped with much cheaper and generalized models like GPT-4o-mini delivers comparable performance, demonstrating its adaptability and scalability. Overall, the Meta-Reasoner framework provides a compatible approach to improving reasoning-intensive tasks, combining high accuracy with dynamic and efficient problem-solving strategies. The results on SciBench and TheoremQA also demonstrate similar findings and show that Meta-Reasoner generally achieves better performance compared to the baselines and the results are consistent across different models.
Model Variant Game-of-24(%) TheoremQA(%)
GPT-4o-mini Full Method 89 84.13
w/o Progress Report 85 79.42
w/o MAB (direct arm selection) 82 80.74
w/o MAB (CoT) 4 39.46
Gemini-Exp-1206 Full Method 94 86.32
w/o Progress Report 91 81.78
w/o MAB (direct arm selection) 87 82.14
w/o MAB (CoT) 11 43.12
Table 5: Ablation study of Meta-Reasoner. Direct arm selection refers to prompting LLM to directly select a strategy based on recent progress report.
Bandit Type Game-of-24(%) #US TheoremQA(%) #US
Fixed (K=3) 65 3 72.34 3
Fixed (K=5) 72 5 79.17 5
Dynamic 89 14 84.13 21
Table 6: Fixed vs. Dynamic Bandit Variants over GPT-4o-mini. #US: Number of Unique Strategies.
Figure 3: The trade-off between accuracy and normalized inference costs across different models and methods on 24-point games. We use gpt-4o-mini as the backend model for all the prompting methods. For each method, key hyper-parameters (e.g., ${N}$ in Best of N, or tree size in ToT) are tuned to yield a baseline (lower point) and an extended (upper point) configuration, with dashed lines connecting these bounds.
Figure 4: Cumulative reward of different settings across iteration. We compare our method using LinUCB with baseline (direct arm selection), and random search methods across two tasks—Game of 24 (top row) and TheoremQA (bottom row) using GPT-40-mini (left) and Gemini-Exp-1206 (right).
5.3 Ablation Study
In this section, we conduct an ablation study to analyze each component contribution of Meta-Reasoner. In specific, we consider the following setup: (1) w/o progress report: we replace the progress reporting process with directly considering the entire CoT history without summarization; (2) w/o MAB: instead of using MAB to select the proper strategy, we directly leverage an LLM to the decision making to provide the proper strategy for LRM reasoning. In Table 5, we show that when removing progress reporting (“w/o Progress Report”), the overall performance moderately degrades and we hypothesize it is due to the concise intermediate summarizations can help the Meta-reasoner only consider the high-level strategy instead of being confused with too much details of the reasoning process. We also find that removing the MAB brings a more pronounced effect, especially when strategy selection falls back to a direct chain-of-thought approach (“w/o MAB (CoT)”). It verifies the effect of our meta-reasoner module to help the model stay on track for getting an optimal solution. In Table 6, we compare fixed and dynamic bandit variants on the game of 24 and theoremQA. We find that using a fixed set of strategies (e.g., $K=3$ and $K=5$) yields lower performance compared to the dynamic approach which adaptively explores more strategies (shown by larger unique strategies). The results highlight the benefit of flexibly allocating diverse reasoning strategies using LLM in-context learning capabilities.
5.4 Analysis
Effectiveness of Meta-reasoner.
In Figure 4, we demonstrate the cumulative rewards across iterations to analyze the effectiveness of the our meta-reasoner module. We compare our MAB-based method with a baseline which directly prompts an LLM to select an arm (“strategy”). We refer to this baseline method as Baseline (Direct Arm Selection), with the prompt in Figure 58. The results show that MAB-based Meta-Reasoner (using LinUCB Li et al. (2012)) consistently outperforms both direct LLM decision-making (the baseline) and random search across two distinct tasks (Game of 24 and TheoremQA) and under two model scales (GPT-40-mini and Gemini-Exp-1206). While direct LLM usage may yield reasonable initial performance and random search incurs minimal setup cost, neither approach systematically balances exploration and exploitation. In contrast, the MAB updating strategy leverages feedback from previous iterations to adaptively refine its action selection (e.g., choosing a proper strategy based on the CoT reasoning), steadily increasing cumulative rewards.
Inference Efficiency. In Figure 3, we compare the inference costs across different models and various prompting strategies. Basic models, like GPT-4o-mini and GPT-4o, show lower accuracy and minimal inference cost, occupying the lower-left corner of the plot. As methods become more advanced, such as ToT and MACM, accuracy improves significantly but at the expense of higher inference costs. Our proposed method stands out by achieving a strong balance between high accuracy and moderate inference cost, outperforming methods like MACM, which delivers lower accuracy at higher costs. While proprietary models like o1-mini and o1-preview achieve slightly higher accuracy, they incur the highest inference costs, highlighting their reliance on more computational resources. Meta-Reasoner demonstrates competitive performance with a cost-effective approach making it a scalable and efficient solution for reasoning-intensive tasks.
6 Conclusion
In this work, we introduce Meta-Reasoner, a meta-reasoning framework designed to enhance the reasoning capabilities of LRMs and optimize the inference-time reasoning efficiency. By operating as an “advisor”, meta-reasoner dynamically evaluates the reasoning process and provides high-level strategic guidance, addressing key limitations of o1-like reasoning chains, such as compounding errors and inefficiency in inference computing. Unlike conventional reasoning approaches, Meta-Reasoner focuses on global oversight rather than granular step-by-step processes, enabling LRMs to avoid unproductive lines of thought and better allocate computational resources. The experiments highlight the potential of dynamic reasoning chains to overcome inherent challenges in the LLM reasoning process and also show promise in broader applications, offering a scalable and adaptable solution for reasoning-intensive tasks.
Limitations
Our proposed Meta-Reasoner framework, while effective at improving inference-time reasoning, has a few key limitations that may affect its applicability. First, it relies on a carefully designed reward function to guide strategy selection: if the reward signal does not accurately reflect correctness or progress, the meta-reasoner may persist with incorrect strategies. This challenge becomes more pronounced when the tasks involve subjective or complex objectives that are hard to quantify automatically (such as creative writing, complex theorem proving). Second, while dynamically adding or refining strategies expands the meta-reasoner’s flexibility, it can also introduce instability. Overly complex or poorly specified new strategies may create confusion rather than enhance problem-solving. Careful vetting or domain-specific constraints may be needed in future version to prevent the system from drifting into ineffective approaches.